OpenStudy (anonymous):

can someone please help me with the formula for this problem the chart for the problem only goes to 10 years.Carla Lopez deposits 3,000.00 in retirement account the funds have an average of 9% over the 40 tears until her retirement what will be the value of her retirement account

OpenStudy (anonymous):

It's not clear to me that the funds are compounding over the 40 years. If we could see the chart that would help. However, if there is compounding over the period the value would be $94,228

OpenStudy (anonymous):

You say that you have a chart that goes to ten years. What is the value at 10 years?

OpenStudy (anonymous):

45579

OpenStudy (anonymous):

i got the answer but I dont cheat I need to find out the formula for the answer

OpenStudy (anonymous):

$45,579 seems big to me unless there is a $3,000 per year contribution over the ten years.

OpenStudy (anonymous):

the chart for 10 years at 9% is 15.193 aren't I suppose to multiply the 3000* the 15.193

OpenStudy (anonymous):

its only 1 deposit of 3000

OpenStudy (anonymous):

Hold on FV = [ 3,000 ( 1.09 ) ^10 - 1] / .09

OpenStudy (anonymous):

what do I do with the 10

OpenStudy (anonymous):

Power 10

OpenStudy (anonymous):

i am so confused

OpenStudy (anonymous):

From 10 years, the future value is $45, 579

OpenStudy (anonymous):

Therefore, all you do is to plug in 40 yrs, instead of 10 yrs.

OpenStudy (anonymous):

is it 3000*1.09 to the 10th power -1 =

OpenStudy (anonymous):

but how

OpenStudy (anonymous):

Just plug into your calculator!

OpenStudy (anonymous):

once i get the answer to the 10th power do I multiply it by 40 years

OpenStudy (anonymous):

No, instead of plug number 10, replace it by 40!

OpenStudy (anonymous):

what do you get for the answer I have 3000*1.09-1 = 3269.0o to the 10th power is to much. we have a financial chart that we go by your correct because I see your example on another problem in the book

OpenStudy (anonymous):

I don't know what calculator you are using, but mine TI 89 shows: $1,013, 650

OpenStudy (anonymous):

in the book its 15.193 for ten years at 9% what do I mulyiply by the 15.193

OpenStudy (anonymous):

For 10 yrs, I got $45,579

OpenStudy (anonymous):

I did to but what do I do with it th 45579

OpenStudy (anonymous):

It just a test to prove which formula to apply!

OpenStudy (anonymous):

I'm still not sure how you came up with your formula. My book says that the future value of a single sum is FVSS=PVSS(1+i)^n.

OpenStudy (anonymous):

@marshallinwashington It's the formula specialized for business application!

OpenStudy (anonymous):

It's the same formula that my HP12c is using. So yes.

OpenStudy (anonymous):

your formula was correct but I don't understand the 40 years do the problem step by step so I can understand I havent been in college for over 20 years just going back FV = [ 3,000 ( 1.09 ) ^10 - 1] / .09 explain this Is fv= 3000*1.09-1 to the 10th power

OpenStudy (anonymous):

In order to calculate the retirement after 40 yrs, power to 40: FV = [ 3,000 ( 1.09 ) ^40 - 1] / .09

OpenStudy (anonymous):

oh ok give me 1 min and i will send you the answer and can you check it please

OpenStudy (anonymous):

Okie dokey!

OpenStudy (anonymous):

do I multiply 1.09 to the 40th power

OpenStudy (anonymous):

power, not multiply!

OpenStudy (anonymous):

like 2^3 = 8

OpenStudy (anonymous):

i thought im suppose to multiply 1.09*40 than multiply the answer by 40 until i do it 40 tmes

OpenStudy (anonymous):

Nope, all I do is enter into my calculator: 3000 ( 1.09^40 - 1) / .09

OpenStudy (anonymous):

Yes, power 40 means repeated multiply 40 times!

OpenStudy (anonymous):

1.09 * 1.09*1.09*1.09*1.09 40 times right

OpenStudy (anonymous):

so the shortcut is power! instead of manually multiply!

OpenStudy (anonymous):

Did you enter what I showed you into your calculator?

OpenStudy (anonymous):

did you get 36925.94

OpenStudy (anonymous):

= $1, 013,647.34

OpenStudy (anonymous):

how

OpenStudy (anonymous):

for 40 yrs

OpenStudy (anonymous):

so once I multiply 1.09 40 times then what do I do

OpenStudy (anonymous):

do I / by .09

OpenStudy (anonymous):

Pls do you know what difference between multiply and power? 2^3 = 8 2 * 3 = 6

OpenStudy (anonymous):

yea 1.09 *1.09 40 times right

OpenStudy (anonymous):

You need to purchase the right calculator, otherwise the result won't be precisely!

OpenStudy (anonymous):

i just pulled up a sciencetific calculator

OpenStudy (anonymous):

So use exponent!

OpenStudy (anonymous):

i got the same answer as you thank you for being very patient with me and god bless you10136.4734

OpenStudy (anonymous):

Just memorize the formula, you'll need it :)

OpenStudy (anonymous):

yea I wrote it down thank you thank you thank you very much

OpenStudy (anonymous):

I guess that's all I know ;)

OpenStudy (anonymous):

You are awesome goodnight

OpenStudy (anonymous):

Anyway, just message me when you need help!

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

Don't forget to click on the blue "Best Answer" as thank you!

OpenStudy (anonymous):

Nett-Nett, Chlorophyl's formula is the future value of an annuity. Which means there was a $3,000 deposit each year. So the high number now makes some sense.